OpenStudy (anonymous):

Which of the following is equal to square root of the cube root of 2? 2 to the power of 2/3 2 to the power of 3/2 2 to the power of 1/3 2 to the power of 1/6 @Jhannybean @TheSmartOne

OpenStudy (jhannybean):

\[\sqrt[n]{x^m}= x^{m/n}\]

OpenStudy (jhannybean):

\[\sqrt[2]{\sqrt[3]{2}} = (\sqrt[3]{2})^{1/2}\] First sep.

OpenStudy (jhannybean):

step*

OpenStudy (anonymous):

2and 3ver 2 is the answer rt

OpenStudy (jhannybean):

Then you take out the 1/3. \[\sf (2^{1/3})^{1/2} = 2^{1/2 \cdot 1/3}\] So what is \(\sf \frac{1}{2} \cdot \frac{1}{3}\)?

OpenStudy (anonymous):

1 over 5

OpenStudy (jhannybean):

@dmndlife24 you read the question wrong.

OpenStudy (anonymous):

n0 6

OpenStudy (jhannybean):

2 x 3 \(\ne\) 5.

OpenStudy (jhannybean):

Good job :)